# On the composable security of weak coin flipping

### v2

@@ Abstract
Weak coin flipping is a cryptographic primitive in which two mutually distrustful parties generate a shared random bit to agree on a winner via remote communication. While a stand-alone secure weak coin flipping protocol can be constructed from noiseless communication channels, its composability has not been explored. In this work, we demonstrate that no weak coin flipping protocol can be abstracted into a simple black box resource with composable security. Despite this, we also establish the overall stand-alone security of weak coin flipping protocols under sequential composition.
@@ 1 Introduction
Several frameworks have been proposed that are intended to analyze compositions of cryptographic systems, starting with Canetti’s seminal Universally Composable (UC) security framework [3, 4]. Since then, several extensions and variants have emerged (see, e.g., [5, 6, 7, 8]). We focus here on the model of Abstract Cryptography (AC) by Maurer and Renner [9, 10]. In contrast to the previous frameworks, the latter is a top-down axiomatic approach and is intended to provide better insights via abstraction of a given complex cryptographic system and will thus be most suitable for our purposes. The security in the AC framework is built on simulation-based security. Here, for a given system to be deemed secure, it suffices to ensure that all adversarial attacks on the system can be simulated on an ideal system. Therefore, to prove the security of any multi-party protocol in the AC framework, one needs to show the existence of such a simulator simulating all possible adversary attacks.
@@
The cryptographic setting we consider in this work is a two-party primitive called weak coin flipping (WCF). Informally speaking, in WCF two mutually distrustful parties, say Alice and Bob, need to agree on a binary outcome even though the two parties have opposite preferences. The protocol is deemed secure if it produces a fair coin toss in case both parties are honest and a dishonest party can at most bias the honest party’s outcome towards their preference. The WCF task is fundamentally different from the other commonly studied tasks such as strong coin flipping (SCF), bit commitment (BC), oblivious transfer, or more generally, secure function evaluation. While the latter protocols are shown to be insecure with an unbounded adversary [11, 12, 13, 14, 15], it is possible to construct weak coin flipping protocols with quantum communication that exhibit almost perfect security with arbitrarily small bias [16]. The explicit protocols which are shown to have an arbitrarily small bias are developed in [17, 18, 19]. The security notion employed in these results, known as information-theoretic security, allows an adversary to enjoy unlimited resources (i.e. the adversary has no restrictions on its computational power or use of memory, etc.). It is well-studied for a range of different cryptographic tasks and, in most cases, is stronger than other commonly studied security definitions such as computational [20, 21, 22], bounded storage [23, 24, 25] or even game-based [26, 27], where an adversary can perform attacks in only some prescribed ways.
@@
Previously, it was shown that WCF enjoys information-theoretic stand-alone security, i.e., it is secure as a single protocol instance in isolation [16, 17, 18, 19]. However, while WCF itself is of limited interest, it is mostly a powerful tool in constructing optimal protocols for several other primitives such as bit commitment [28] and strong coin flipping [29]. However, such constructions would require more than stand-alone security and it is still unclear if WCF protocols have composable security when combined with other cryptographic protocols.
@@
Our main contributions encompass two negative theorems and one positive theorem. In Theorem 5, we demonstrate the impossibility of constructing a typical black-box WCF resource analogous to the widely-used SCF resource [30] (see Definition 4) from noiseless communication channels in the AC framework. This WCF resource generates a random bit for both parties, allowing the cheating party to learn the bit in advance and bias it towards their preferred direction. It was shown that the widely-used SCF resource cannot be achieved in the AC framework [31], but the result for the similar WCF resource remains unknown until this work. We state Theorem 5 here informally.
@@
The typical black-box WCF resource with composable security cannot be constructed from noiseless communication channels in the AC framework.
To prove Theorem 5, we show that WCF protocols using noiseless communication channels are vulnerable to the man-in-the-middle attack when composed in parallel. As a result, the typical black-box WCF resource cannot be constructed by any WCF protocol from noiseless communication channels in the AC framework.
Theorem 5 shows the impossibility of the typical black-box WCF resource in the AC framework. However, we can still define an alternative WCF resource at the cost of increasing the interaction complexity. Nevertheless, such an alternative WCF resource that can be constructed from noiseless communication channels must be extremely complicated in terms of its interaction complexity, as is shown in Theorem 6. Replacing an actual protocol with an ideal resource intends to simplify the analysis; therefore, it is meaningless to study such an alternative WCF resource. Here is the informal statement of Theorem 6.
@@
Any black-box WCF resource with composable security constructed from noiseless communication channels in the AC framework must have high interaction complexity.
The proof of Theorem 6 consists of two steps. First, we show that an alternative WCF resource with composable security constructed from noiseless communication channels implies a WCF protocol with the same interaction complexity as the alternative WCF resource. Then together with the exponential lower bound on the interaction complexity of WCF protocols, we conclude that a WCF resource with a low interaction complexity can not exist.
Next, we extend the classical version of the sequential composition theorem [32, Theorem 7.4.3] to the quantum case. We find a sufficient condition called global security where WCF and other modules can be composed into a larger system with stand-alone security. A composition between WCF and other components is globally secure if there is no correlation between the outcomes of WCF and other components. We thus show in Theorem 10 that global security holds when WCF and all other components are composed in a strict causal order. That is, there is no other component running concurrently with WCF. The constructions for optimal SCF and optimal BC exactly satisfy global security.
@@
A primitive constructed from WCF with stand-alone security under strict sequential composition maintains its stand-alone security.
The idea to prove Theorem 10 is by contradiction. Assume that there is some correlation between outcomes of WCF with stand-alone security and other components under sequential composition. Using the information carried from previous components to WCF, we can construct an attack breaking the stand-alone security of WCF. This contradicts our assumption and thus proves Theorem 10.
@@
Causality in the composable framework. Some early works [33, 34] developed the paradigm that only allows for sequential composition, where protocols are executed one after another. Canetti [3] relaxed “sequential" composition to modular composition in the UC framework, where the former is a special case of the latter. The AC framework [9, 10] takes a top-down hierarchy. Only the connection of abstract systems is considered at the top level, and the execution order is left at the lower level specification. Under the AC framework, the causal box model [35] provides the causal order specification that allows for quantum superposition of different orders.
Concerning concurrent composition. Literature [24, 25] addresses the difficulty of concurrent parallel composition in the UC framework and the bounded storage model. The composability of BC was discussed in [36]. They define the weak binding security of BC in a stand-alone rather than composable manner and propose a counterexample where the string commitment constructed from multiple instances of BC is insecure. In a previous model of general cryptographic protocols [32, Section 7.3], concurrent calls to an oracle (resource) are also forbidden. This restriction primarily stems from the formulation of syntax in that model rather than an explicit security concern.
@@ 2.1 The abstract cryptography framework
The building blocks of the framework are resources, converters (protocols), and distinguishers. Resources are cryptographic functionalities, and converters transform one resource into another. Distinguishers distinguish resources from each other. The framework is feasible to describe multiparty functionalities, but we only focus on two-party ones here.
  • A resource, denoted by $R$, has one interface accessible to each party. Each party can send inputs, receive outputs, perform operations and give instructions via the corresponding interface. In two-party cryptography, it is necessary to consider three different behaviours $R$, $R_{A}$ and $R_{B}$ of a resource, where $R$, $R_{A}$ and $R_{B}$ correspond to the case where both are honest, where Alice is dishonest, and where Bob is dishonest, respectively. We use $\mathcal{R}=(R,R_{A},R_{B})$ to denote a resource compactly.
  • A converter, denoted by a lower-case Greek letter $\alpha$, is a cryptographic system with an inner interface connected to an interface of a resource and an outer interface to a party. The set of all converters is denoted by $\Sigma$. Each party can instruct the converter on the outer interface to execute specific tasks on the inner interface. A converter $\alpha$ can convert one resource $R$ into another resource $S$ by connecting its inner interface to one interface of the resource $R$. The interfaces of the resource $S$ consist of the unconnected interface of the resource $R$ and the outer interface of the converter $\alpha$. We denote the connection between the converter $\alpha$ and the left-side (right-side) interface of the resource $R$ by $\alpha R$ ($R\alpha$).
  • A protocol $\pi=(\pi_{A},\pi_{B})$ consists of Alice’s converter $\pi_{A}$ and Bob’s converter $\pi_{B}$. $\pi_{A}$ and $\pi_{B}$ are operations that honest Alice and Bob should follow when building one resource from another. By connecting each converter to the corresponding interface, the protocol $\pi$ can build a required resource $S$ from a given resource $R$ if $S=\pi_{A}R\pi_{B}$. For example, suppose $R$ is a binary symmetric channel resource and $(\pi_{A}$,$\pi_{B})$ is an encoder-decoder pair corresponding to a proper error-correcting code. Then one can construct a noiseless channel resource as $S=\pi_{A}R\pi_{B}$.
  • A distinguisher, denoted by $\mathcal{D}$, is an agent who can access both interfaces of a resource. Given one of two resources $R$ and $S$ uniformly at random, a distinguisher guesses whether the resource is $R$ or $S$ by interacting with it. The distinguishing advantage of the above task is defined as
@@
    The distinguishing advantage is a pseudo metric between two resources, satisfying the symmetry and triangle inequality properties. When $d(R,S)\leq\delta$, we also write $R\approx_{\delta}S$.
@@
A protocol $\pi=(\pi_{A},\pi_{B})$ constructs resource $\mathcal{S}=(S,S_{A},S_{B})$ out of resource $\mathcal{R}=(R,R_{A},R_{B})$ with distinguishing advantage $\delta$ if the following conditions are satisfied
@@
    $\exists\sigma_{B}\in\Sigma,S_{B}\sigma_{B}\approx_{\delta}\pi_{A}R_{B}.$ (3)
@@
In the previous definition, $\sigma_{A}$ and $\sigma_{B}$ are called simulators. Intuitively, the simulators here ensure that any valid cheating strategy for resource $\mathcal{R}$ could be mapped onto some cheating strategy for resource $\mathcal{S}$. For instance, if dishonest Bob deviates from the protocol $\pi_{B}$ by following $\pi^{\prime}_{B}$ instead, then, according to Eq. (3), this is equivalent to the action $\sigma_{B}\pi^{\prime}_{B}$ on resource $S_{B}$. For a detailed description of simulators and the related security notions in cryptography, we refer to the tutorial [37].
@@ 2.2 The quantum comb model
A quantum comb [38, 39] is a general quantum information processing system with internal memory that processes the incoming messages one by one. An example of quantum comb is shown in Fig. 1. It is common to model the building blocks in the AC framework as quantum comb. If a protocol can be modeled with the quantum comb model (QCM), it is called a QCM protocol. Note that all existing WCF protocols can be classified as QCM protocols.
@@
Figure 1: An example of quantum comb. $\mathcal{E}_{i}$ is a quantum operation. The lines with an arrow are quantum states. The unconnected lines are pairs of input and output messages.
The interaction complexity of a QCM protocol $\pi$, denoted by $\mathsf{Intcom}(\pi)$, is defined as its communication rounds. As an example, for a protocol $\pi$ modeled by Fig. 1, $\mathsf{Intcom}(\pi)=n$. For resources such as $S_{A}$ and $S_{B}$, their interaction complexity refers to the communication rounds at the adversary’s side. For example, the interaction complexity of WCF resource $S_{A}$ in Fig 3(b) is the rounds at the left side, which is $1$.
@@ 2.3 The causal box model
The causal box model [35] is a more general specification of building blocks in the AC framework. The causal box model presents the causal structure of the messages sent or received by these building blocks. The quantum comb model can be treated as a special case of the causal box model.
In the causal box model, the abstract blocks are modeled as causal boxes. A causal box is an information processing system with input and output wires. Each message on a wire is described by a quantum state on space $\mathcal{H}\otimes l^{2}(\mathcal{T})$, where $\mathcal{T}$ is a partially ordered set (poset) corresponding to the order of messages and $l^{2}(\mathcal{T})$ is the set of all functions $\mathcal{T}\to\mathbb{C}$ with bounded $L^{2}$ norm. Note that $\mathcal{T}$ cannot be simply interpreted as time, which is totally ordered. For generality, a wire may contain arbitrarily many messages. Therefore, the wire space is defined as the Fock space of the message space
@@
where $\vee^{n}\mathcal{H}$ denotes the symmetric subspace of $\mathcal{H}^{\otimes n}$ and $l^{2}(\mathcal{T})$ is the space of square summable sequence on $\mathcal{T}$. Sometimes we abuse the notation by writing $\mathcal{F}_{\mathcal{H}_{A}}^{\mathcal{T}}$ as $\mathcal{F}_{A}^{\mathcal{T}}$. With this definition, wires can be merged or split, i.e., for $\mathcal{H}_{A}=\mathcal{H}_{A_{1}}\oplus\mathcal{H}_{A_{2}}$,
@@
A causal box is formally defined as a completely positive and trace-preserving (CPTP) map from the input wire $X$ to the output wire $X^{\prime}$. This map should further respect the causality condition, i.e., the output message at $t^{\prime}\in\mathcal{T}$ only depends on the input messages at $t\in\mathcal{T}$ where $t\preceq t^{\prime}$. A precise definition can be found in Appendix A.
The composition of causal boxes consists of parallel composition and loops. The parallel composition of causal boxes $\Phi$ and $\Psi$, $\Phi\|\Psi$ is defined as the map $\Phi\otimes\Psi$. Due to Equation (9), this map’s input and output space are still valid wires. A loop connects segments with the same dimension of the input and output wires. An interface labelled by $x$ of a resource or a converter $\Phi$, denoted by $\mathsf{int}_{x}(\Phi)$, is specified by a set of wires, possibly including both input and output wires. For example, the inner and outer interfaces of a converter $\alpha$ are denoted by $\mathsf{int}_{in}(\alpha)$ and $\mathsf{int}_{out}(\alpha)$, respectively. Two interfaces $\mathsf{int}_{a}(\Phi),\mathsf{int}_{b}(\Psi)$ are compatible if there exists a bijection $P:\mathsf{int}_{a}(\Phi)\leftrightarrow\mathsf{int}_{b}(\Psi)$ such that each pair of wires consists of an input wire and an output wire with the same dimension.

### v3

@@ Abstract
Weak coin flipping is a cryptographic primitive in which two mutually distrustful parties generate a shared random bit to agree on a winner via remote communication. While a stand-alone secure weak coin flipping protocol can be constructed from noiseless quantum communication channels, its composability remains unexplored. In this work, we demonstrate that no weak coin flipping protocol can be abstracted as a simple black-box resource with composable security. Despite this, we also establish the overall stand-alone security of quantum weak coin flipping protocols under composition in sequential order.
@@ 1 Introduction
Several frameworks have been proposed that are intended to analyze compositions of cryptographic systems, starting with Canetti’s seminal Universally Composable (UC) security framework [3, 4]. Since then, several extensions and variants have emerged (see, e.g., [5, 6, 7, 8]). We focus here on the model of Abstract Cryptography (AC) by Maurer and Renner [9, 10]. In contrast to the previous frameworks, the latter is a top-down axiomatic approach and is intended to provide better insights via abstraction of a given complex cryptographic system and is thus most suitable for our purposes. The security in the AC framework is built on simulation-based security. Here, for a given system to be deemed secure, it suffices to ensure that all adversarial attacks on the system can be simulated on an ideal system. Therefore, to prove the security of any multi-party protocol in the AC framework, one needs to show the existence of such a simulator simulating all possible adversary attacks.
@@
The cryptographic setting we consider in this work is a two-party primitive called weak coin flipping (WCF). Informally speaking, in WCF two mutually distrustful parties, say Alice and Bob, need to agree on a binary outcome even though the two parties have opposite preferences. The protocol is deemed secure if it produces a fair coin toss in case both parties are honest and a dishonest party can at most bias the honest party’s outcome towards their preference. The WCF task is fundamentally different from the other commonly studied tasks such as strong coin flipping (SCF), bit commitment (BC), oblivious transfer, or more generally, secure function evaluation. While the latter protocols are shown to be insecure with an unbounded adversary [11, 12, 13, 14, 15], it is possible to construct weak coin flipping protocols with quantum communication that exhibit almost perfect security with arbitrarily small bias [16]. The explicit protocols which are shown to have an arbitrarily small bias are developed in [17, 18, 19]. The security notion employed in these results, known as information-theoretic security, allows an adversary to enjoy unlimited resources (i.e., the adversary has no restrictions on its computational power or use of memory, etc.). It is well-studied for a range of different cryptographic tasks and, in most cases, is stronger than other commonly studied security definitions such as computational [20, 21, 22], bounded storage [23, 24, 25] or even game-based [26, 27], where an adversary can perform attacks in only some prescribed ways.
@@
Previously, it was shown that WCF enjoys information-theoretic stand-alone security, i.e., it is secure as a single protocol instance in isolation [16, 17, 18, 19]. While WCF itself is of limited interest, it is a powerful tool in constructing optimal protocols for several other primitives such as bit commitment [28] and strong coin flipping [29]. However, such constructions would require more than stand-alone security, and it is still unclear if WCF protocols have composable security when combined with other cryptographic protocols.
@@
Our main contributions encompass two negative theorems and one positive theorem. In Theorem 5, we demonstrate the impossibility of constructing a typical WCF resource (see Definition 4) analogous to the widely-used SCF resource [30] from noiseless communication channels in the AC framework. Generally speaking, a resource is the abstraction of certain cryptographic functionality (see Section 2.1). This WCF resource generates a random bit for both parties, allowing the cheating party to learn the bit in advance and bias it towards their preferred choice. It was shown that the widely-used SCF resource cannot be achieved in the AC framework [31], but the result for the similar WCF resource remains unknown until this work. We state Theorem 5 here informally.
@@
The typical WCF resource in Definition 4 with composable security cannot be constructed from noiseless communication channels in the AC framework.
To prove Theorem 5, we show that WCF protocols using noiseless communication channels are vulnerable to the man-in-the-middle attack when composed in parallel. As a result, the typical WCF resource cannot be constructed by any WCF protocol from noiseless communication channels in the AC framework.
While Theorem 5 shows the impossibility of the typical WCF resource in the AC framework, we can still define an alternative WCF resource at the cost of increasing the interaction complexity. Nevertheless, such an alternative WCF resource that can be constructed from noiseless communication channels must be extremely complicated in terms of its interaction complexity, as is shown in Theorem 6. Replacing an actual protocol with an ideal resource intends to simplify the analysis; therefore, it is meaningless to study such an alternative WCF resource. Here is the informal statement of Theorem 6.
@@
Any WCF resource with composable security constructed from noiseless communication channels in the AC framework must have high interaction complexity.
The proof of Theorem 6 consists of two steps. First, we show that an alternative WCF resource with composable security constructed from noiseless communication channels implies a WCF protocol with the same interaction complexity as the alternative WCF resource. Then, together with the exponential lower bound on the interaction complexity of WCF protocols [32], we conclude that a WCF resource with a low interaction complexity cannot exist.
Next, we extend the classical version of the composition theorem for sequentially ordered protocols [33, Theorem 7.4.3] to the quantum case. We find a sufficient condition called global security where WCF and other modules can be composed into a larger system with stand-alone security. A composition between WCF and other components is globally secure if there is no correlation between the outcomes of WCF and other components. We thus show in Theorem 10 that global security holds when WCF and all other components are composed in a strict causal order. That is, there is no other component running concurrently with WCF. The constructions for optimal SCF and optimal BC exactly satisfy global security.
@@
A primitive constructed from WCF with stand-alone security under strictly ordered sequential composition maintains its stand-alone security.
The idea to prove Theorem 10 is by contradiction. Assume that there is some correlation between outcomes of WCF with stand-alone security and other components under ordered sequential composition. Using the information carried from previous components to WCF, we can construct an attack breaking the stand-alone security of WCF. This contradicts our assumption and thus proves Theorem 10.
@@
Causality in the composable framework. Some early works [34, 35] developed the paradigm that only allows for sequential-order composition, where protocols are executed one after another. Canetti [3] relaxed sequential composition to modular composition in the UC framework, where the former is a special case of the latter. The AC framework [9, 10] takes a top-down approach. Only the connection of abstract systems is considered at the top level, and the execution order is left at the lower level specification. Under the AC framework, the causal box model [36] provides the causal order specification that allows for a quantum superposition of different message orders.
Concerning concurrent composition. Literature [24, 25] addresses the difficulty of concurrent parallel composition in the UC framework and the bounded-storage model. The composability of BC was discussed in [37]. They define the weak binding security of BC in a stand-alone rather than composable manner and propose a counterexample where the string commitment constructed from multiple instances of BC is insecure. In a previous model of general cryptographic protocols [33, Section 7.3], concurrent calls to an oracle (resource) are also forbidden. This restriction primarily stems from the formulation of syntax in that model, rather than an explicit security concern. Recent work by Batra et al. [38] provides device-independent protocols for oblivious transfer and bit commitment proven secure under sequential composition, where all queries to the ideal functionality must be performed sequentially, but their concurrent composable security remains unknown.
@@ 2.1 The abstract cryptography framework
The building blocks of the framework are resources, converters (protocols), and distinguishers. Resources are interactive systems that abstract cryptographic functionalities, and converters transform one resource into another. Distinguishers distinguish resources from each other. The framework can describe multiparty functionalities, but in this work, we only focus on two-party ones.
  • A resource, denoted by $R$, has one interface accessible to each party. Each party can send inputs, receive outputs, perform operations and give instructions via the corresponding interface. In two-party cryptography, it is necessary to consider three different behaviors $R$, $R_{A}$ and $R_{B}$ of a resource, in which $R$, $R_{A}$ and $R_{B}$ correspond to the case where both are honest, where only Alice is dishonest, and where only Bob is dishonest, respectively. We use $\mathcal{R}=(R,R_{A},R_{B})$ to compactly denote a resource.
  • A converter, denoted by $\alpha$, is a cryptographic system with an inner interface that connects to a resource and an outer interface that connects to a player. The set of all converters is denoted by $\Sigma$. Each party can instruct the converter on the outer interface to execute specific tasks on the inner interface. A converter $\alpha$ can convert one resource $R$ into another resource $S$ by connecting its inner interface to one interface of the resource $R$. The interfaces of the resource $S$ consist of the unconnected interface of the resource $R$ and the outer interface of the converter $\alpha$. We denote the connection between the converter $\alpha$ and the left-side (right-side) interface of the resource $R$ by $\alpha R$ ($R\alpha$).
  • A protocol $\pi=(\pi_{A},\pi_{B})$ consists of Alice’s converter $\pi_{A}$ and Bob’s converter $\pi_{B}$. By connecting each converter to the corresponding interface of a resource $R$, the protocol $\pi$ can transform $R$ into the required resource $S$, provided that $S=\pi_{A}R\pi_{B}$. For example, suppose $R$ is a binary symmetric channel resource and $(\pi_{A}$,$\pi_{B})$ is an encoder-decoder pair corresponding to a proper error-correcting code. Then one can construct a noiseless channel resource as $S=\pi_{A}R\pi_{B}$.
  • A distinguisher, denoted by $\mathcal{D}$, is an agent who can access both interfaces of a resource. Given one of two resources $R$ and $S$ uniformly at random, a distinguisher distinguishes whether the resource is $R$ or $S$ by interacting with it on both interfaces. The distinguishing advantage of the above task is defined as
@@
    The distinguishing advantage is a pseudo metric between two resources, satisfying the symmetry and triangle inequality properties. If $d(R,S)\leq\delta$, we also write $R\approx_{\delta}S$.
@@
A protocol $\pi=(\pi_{A},\pi_{B})$ constructs resource $\mathcal{S}=(S,S_{A},S_{B})$ out of resource $\mathcal{R}=(R,R_{A},R_{B})$ with distinguishing advantage $\delta$ if the following conditions are satisfied:
@@
    $\exists\sigma_{B}\in\Sigma,S_{B}\sigma_{B}\approx_{\delta}\pi_{A}R_{B};$ (3)
@@
In Definition 1, $\sigma_{A}$ and $\sigma_{B}$ are called simulators. These simulators enforce a critical security property: any cheating strategy against the real resource $\mathcal{R}$ can be translated via the simulators into a corresponding cheating strategy against the ideal resource $\mathcal{S}$. For instance, if dishonest Bob deviates from $\pi_{B}$ and follows $\pi^{\prime}_{B}$ instead, then, we have $S_{B}\sigma_{B}\pi^{\prime}_{B}\approx_{\delta}\pi_{A}R_{B}\pi_{B}^{\prime}$ by Equation 3, which means Bob’s strategy $\pi_{B}^{\prime}$ in the real world (interacting with $\pi_{A}R_{B}$) is simulated by the $\sigma_{B}\pi^{\prime}_{B}$ in the ideal world (interacting with $S_{B}$). For a detailed description of simulators and the related security notions in cryptography, we refer to the tutorial [39].
@@ 2.2 The quantum comb model
A quantum comb [40, 41] is a general quantum information processing system with internal memory that processes incoming messages sequentially. An example of a quantum comb is shown in Figure 1. At the lower level specification of the AC framework where the execution order is taken into account, building blocks are commonly modeled as quantum combs. A protocol which can be modeled within the quantum comb model (QCM) is called a QCM protocol. All existing WCF protocols can be classified as QCM protocols.
@@
Figure 1: An example of quantum comb. The lines represent quantum registers, and $\mathcal{E}_{i}$ is a quantum operation. Each quantum operation interacts with the environment through an input-output pair, resembling the teeth of a comb.
The interaction complexity of a QCM protocol $\pi$, denoted by $\mathsf{Intcom}(\pi)$, is defined as the communication rounds it requires. As an example, for a protocol $\pi$ modeled by Figure 1, $\mathsf{Intcom}(\pi)=n$. For resources $S_{A}$ and $S_{B}$, their interaction complexity refers to the communication rounds at the adversary’s side. For example, the interaction complexity of WCF resource $S_{A}$ in Figure 3(b) is the number of rounds on the left side, which is $1$.
@@ 2.3 The causal box model
The causal box model [36] is a more general specification of building blocks in the AC framework. The causal box model describes the causal structure of the messages exchanged by these building blocks. The quantum comb model can be treated as a special case of the causal box model.
In the causal box model, the abstract blocks are modeled as causal boxes. A causal box is an information processing system with input and output wires. Each message on a wire is described by a quantum state on space $\mathcal{H}\otimes l^{2}(\mathcal{T})$, where $\mathcal{T}$ is a partially ordered set (poset) that defines the message ordering, and $l^{2}(\mathcal{T})$ is the set of all functions $\mathcal{T}\to\mathbb{C}$ with bounded $L^{2}$-norm. Note that $\mathcal{T}$ is not necessarily interpreted as time, as time is totally ordered. In general, a wire may carry multiple messages. Therefore, the wire space is defined as the Fock space of the message space,
@@
where $\vee^{n}\mathcal{H}$ denotes the symmetric subspace of $\mathcal{H}^{\otimes n}$. Here, $\mathcal{F}_{\mathcal{H}}^{\mathcal{T}}$ is also a Hilbert space. For convenience, we use the shorthand notation $\mathcal{F}_{A}^{\mathcal{T}}$ for $\mathcal{F}_{\mathcal{H}_{A}}^{\mathcal{T}}$. With this definition, wires can be merged or split, i.e., for $\mathcal{H}_{A}=\mathcal{H}_{A_{1}}\oplus\mathcal{H}_{A_{2}}$,
@@
Formally, a causal box is defined as a completely positive and trace-preserving (CPTP) map that transforms the input wires to the output wires. This map should further respect the causality condition, i.e., the output message at $t^{\prime}\in\mathcal{T}$ only depends on the input messages at $t\in\mathcal{T}$ where $t\preceq t^{\prime}$. A precise definition can be found in Appendix A.
The composition of causal boxes consists of parallel composition and loops. The parallel composition of causal boxes $\Phi$ and $\Psi$, denoted by $\Phi\|\Psi$, is defined as the map $\Phi\otimes\Psi$. By Equation 9, the input and output spaces of $\Phi\otimes\Psi$ remain valid wires. A loop connects segments with the same dimension of the input and output wires. An interface labeled by $x$ for a causal box $\Phi$ is denoted by $\mathsf{int}_{x}(\Phi)$, which consists of a set of wires, possibly including both input and output wires. For example, the inner and outer interfaces of a converter $\alpha$ are denoted by $\mathsf{int}_{in}(\alpha)$ and $\mathsf{int}_{out}(\alpha)$, respectively. Two interfaces $\mathsf{int}_{a}(\Phi)$ and $\mathsf{int}_{b}(\Psi)$ are compatible if there exists a bijection $P:\mathsf{int}_{a}(\Phi)\leftrightarrow\mathsf{int}_{b}(\Psi)$ such that each input wire of $\mathsf{int}_{a}(\Phi)$ corresponds to an output wire of $\mathsf{int}_{b}(\Psi)$ with the same dimension and vice versa.